# Photoinduced melting dynamics and collective mode in a correlated charge-order system

### v1

@@ II.3 Many-body dynamics with exact diagonalization
and diagonalize ${\mathcal{H}}_{\rm tot}={\mathcal{H}}+{\mathcal{H}}_{\rm sp}$ using the Lanczos method. The time evolution of the system under pulsed laser light is simulated by numerically solving the time-dependent Schrödinger equation [35, 36, 38, 39, 40, 37],
@@ III.1 Time-dependent Hartree-Fock approximation
with $A_{\rm cw}=0.01$ and $\gamma=0.02$, and compute the increase in total energy at sufficiently large $\tau$. Because of the existence of the collective mode, $\alpha(\omega)$ has a sharp peak at $\omega=\Omega_{c}=2.3$ ($\simeq\Delta_{\rm CO}/2$[31]. For $\omega>\Delta_{\rm CO}$, a broad absorption band appears, which is due to interband single-particle excitations.
@@ III.2 Many-body dynamics with exact diagonalization
In this section, we show results of the calculations using the ED method. We should note that, compared to the time-dependent HF method discussed above, quantum fluctuations can be fully treated but the cluster size one can consider due to numerical limitation is considerably smaller so that finite-size effects should be carefully characterized. The charge gap before light irradiation is defined as $\Delta_{\rm CO}=\mu_{+}-\mu_{-}$ with $\mu_{+}=(E_{0}(N,N_{e}+2)-E_{0}(N,N_{e}))/2$ and $\mu_{-}=(E_{0}(N,N_{e})-E_{0}(N,N_{e}-2))/2$ where $E_{0}(N,N_{e})$ is the ground state energy for a system of $N$ sites with $N_{e}$ electrons. For $V=3$, we obtain $\Delta_{\rm CO}=2.42$ for $N=22$ and $\Delta_{\rm CO}=2.79$ for $N=18$.
Figure 5(a) shows the linear optical absorption spectra $\alpha(\omega)$ calculated from the many-body dynamics for the two cluster sizes, where we use $A(\tau)$ in Eq. (13) with $A_{\rm cw}=0.01$ and $\gamma=0.02$. The spectra consist from discrete peaks, which is due to the finite size effect [49, 50]. The peak at $\omega=\Omega_{c}=1.0$ (which is slightly lower than $\Delta_{\rm CO}/2$, but is actually half of the optical gap as we will see later) corresponds to the collective phase mode of the CO that has the lowest excitation energy of the system, as in the case of HF approximation shown in Fig. 1(a). In fact, by calculating up to the 3rd lowest eigenvalue, $E_{2}(N,N_{e})$, we obtain the lowest excitation energy from doubly-degenerate ground states as $E_{2}(N,N_{e})-E_{0}(N,N_{e})=0.94$ for $N=22$, which is consistent with the peak position $\omega=\Omega_{c}$ in $\alpha(\omega)$ and thus supports our assignment. We note that other peaks with $\omega>\Omega_{c}$ in $\alpha(\omega)$ might come from excitation of collective motions of CO with nonzero momenta [49]. The finite size effect for the frequency $\Omega_{c}$ is small, whereas the other peak frequencies vary largely with $N$.
@@
In Figs. 6(a) and 6(b), we show $D^{<}(\varepsilon,\tau_{\rm pr})$ and $D(\varepsilon,\tau_{\rm pr})$ with $\omega_{p}=1.0$, respectively, for $\tau_{\rm pr}=10$ and $N=18$. Without the light ($A_{p}=0$), $D(\varepsilon,\tau_{\rm pr})$ indicates an optical gap of about 2.0 [Fig. 6(b)], which is smaller than $\Delta_{\rm CO}$. For $A_{p}\neq 0$, there appears peak-like in-gap structure at $\varepsilon\simeq 0$ whose intensity increases with increasing $A_{p}$. For $A_{p}=2.4$, the gap in $D(\varepsilon,\tau_{\rm pr})$ is collapsed from inside the gap structure as shown in Fig. 6(b). This characteristic behavior is qualitatively consistent with those obtained by the HF approximation shown in Fig. 3. The notable difference between the ED and HF results is that spectral weights at low energies ($4.0\lesssim\varepsilon\lesssim-3.0$) in $D^{<}(\varepsilon,\tau_{\rm pr})$ decrease more largely in the ED results, which may be due to the correlation effect [51, 52].
@@
Figure 8: Momentum dependence of $A^{<}_{k}(\varepsilon,\tau_{\rm pr})$ for (a) $A_{p}=0$, (b) $\omega_{p}=1.0$ and $A_{p}=2.4$, and (c) $\omega_{p}=5.0$ and $A_{p}=2.4$. The dotted line indicates $\varepsilon=0$. Other parameters are $\tau_{\rm pr}=10$, $\tau_{p}=3$, and $N=18$.
To examine the results in more detail, we show the momentum dependence of $A_{k}^{<}(\varepsilon,\tau_{\rm pr})$ for different $\omega_{p}$ and $A_{p}$ in Fig. 8. For $A_{p}=0$ (no photo-irradiation), a large peak at each $k$ forms the valence band of the CO ground state [Fig. 8(a)]. For $\omega_{p}=1.0$ and $A_{p}=2.4$, the in-gap weights exist at $k=4\pi/9$, $5\pi/9$ ($\simeq\pi/2$) which are connected to peaks forming a dispersion similar to the valence band, shifted to higher energy, as shown in Fig. 8(b). This is consistent with the HF results for $\omega_{p}\simeq\Omega_{c}$ in Figs. 2(a) and 2(b). For $\omega_{p}=5.0$ and $A_{p}=2.4$, spectral peaks in the ground state decrease and there is a spectral transfer to $\varepsilon=1.8$ at $k=2\pi/9$, $7\pi/9$ without notable dispersion around them, as shown in Fig. 8(c). This indicates the interband excitations as in the HF results shown in Fig. 2(d).
@@
Finally, we discuss the decrease in $\phi$ due to photoexcitation by varying $\omega_{p}$ and $A_{p}$ [31, 32]. In Fig. 9, we show the two-dimensional plot of $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane for $\tau_{p}=3$ and $N=22$, where $\phi_{0}$ and $\phi_{\rm av}$ are the CO order parameter in the ground state and the time-averaged $\phi$ in the time domain of $10<\tau<30$, respectively. The structure of the two-dimensional plot can be understood essentially from the behavior of $\alpha(\omega)$ shown in Fig. 5(a); for $\omega_{p}$ with large $\alpha(\omega)$, $\phi$ efficiently decreases with small $A_{p}$, whereas for $\omega_{p}$ with small $\alpha(\omega)$, a large $A_{p}$ is needed to suppress $\phi$. In the HF results [31, 32], the decrease in $\phi$ is basically small when $\omega_{p}$ is not resonant with the collective mode. However, this does not hold in the ED results, which may be due to photoinduced collective motions of CO with nonzero momenta [49] that are not captured in the HF approximation.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Number 23K03309.

### v2

# Photoinduced melting dynamics and collective mode in a correlated charge-ordered system

@@ II.3 Many-body dynamics with exact diagonalization
and diagonalize ${\mathcal{H}}_{\rm tot}={\mathcal{H}}+{\mathcal{H}}_{\rm sp}$ using the Lanczos method where the periodic boundary condition is applied. We consider systems with $N=18$ and $22$ for which the closed shell is realized at half-filling. The time evolution of the system under pulsed laser light is simulated by numerically solving the time-dependent Schrödinger equation [35, 36, 38, 39, 40, 37],
@@ III.1 Time-dependent Hartree-Fock approximation
with $A_{\rm cw}=0.01$ and $\gamma=0.02$, and compute the increase in total energy at sufficiently large $\tau$. We note that $\alpha(\omega)$ is directly related to the regular part of optical conductivity $\sigma_{\rm reg}(\omega)$, which we discuss in Appendix A. Because of the existence of the collective mode, $\alpha(\omega)$ has a sharp peak at $\omega=\Omega_{c}=2.3$ ($\simeq\Delta_{\rm CO}/2$[31]. For $\omega>\Delta_{\rm CO}$, a broad absorption band appears, which is due to interband single-particle excitations.
@@ III.2 Many-body dynamics with exact diagonalization
In this section, we show results of the calculations using the ED method. We should note that, compared to the time-dependent HF method discussed above, quantum fluctuations can be fully treated but the cluster size one can consider due to numerical limitation is considerably smaller so that finite-size effects should be carefully characterized. The charge gap before light irradiation is defined as $\Delta_{\rm CO}=\mu_{+}-\mu_{-}$ with $\mu_{+}=E_{0}(N,N_{e}+1)-E_{0}(N,N_{e})$ and $\mu_{-}=E_{0}(N,N_{e})-E_{0}(N,N_{e}-1)$ where $E_{0}(N,N_{e})$ is the ground state energy for a system of $N$ sites with $N_{e}$ electrons. For $V=3$ and at half filling $(N_{e}=N/2)$, we obtain $\Delta_{\rm CO}=1.77$ for $N=22$ and $\Delta_{\rm CO}=1.99$ for $N=18$.
Figure 5(a) shows the linear optical absorption spectra $\alpha(\omega)$ calculated from the many-body dynamics for the two cluster sizes, where we use $A(\tau)$ in Eq. (13) with $A_{\rm cw}=0.01$ and $\gamma=0.02$. The spectra consist from discrete peaks, which is due to the finite size effect [49, 50]. The peak at $\omega=\Omega_{c}=1.0$ (which is slightly larger than $\Delta_{\rm CO}/2$, but is actually half of the single-particle gap in $D(\varepsilon,\tau_{\rm pr}$) as we will see later) corresponds to the collective phase mode of the CO that has the lowest excitation energy of the system, as in the case of HF approximation shown in Fig. 1(a). In fact, by calculating up to the 3rd lowest eigenvalue, $E_{2}(N,N_{e})$, we obtain the lowest excitation energy from doubly-degenerate ground states as $E_{2}(N,N_{e})-E_{0}(N,N_{e})=0.94$ for $N=22$ and 1.07 for $N=18$, which is consistent with the peak position $\omega=\Omega_{c}$ in $\alpha(\omega)$ and thus supports our assignment. We note that other peaks with $\omega>\Omega_{c}$ in $\alpha(\omega)$ might come from excitation of collective motions of CO with nonzero momenta [49]. The finite size effect for the frequency $\Omega_{c}$ is small, whereas the other peak frequencies vary largely with $N$.
@@
In Figs. 6(a) and 6(b), we show $D^{<}(\varepsilon,\tau_{\rm pr})$ and $D(\varepsilon,\tau_{\rm pr})$ with $\omega_{p}=1.0$, respectively, for $\tau_{\rm pr}=10$. In the calculations of transient spectra, we use $N=18$ due to computational constraint. We note that the results for $N=14$ are qualitatively consistent with those for $N=18$ and therefore the finite size effect does not alter our discussions below. Without the light ($A_{p}=0$), $D(\varepsilon,\tau_{\rm pr})$ indicates a single-particle gap of about 2.0 [Fig. 6(b)], which is close to $\Delta_{\rm CO}$. For $A_{p}\neq 0$, there appears peak-like in-gap structure at $\varepsilon\simeq 0$ whose intensity increases with increasing $A_{p}$. For $A_{p}=2.4$, the gap in $D(\varepsilon,\tau_{\rm pr})$ is collapsed from inside the gap structure as shown in Fig. 6(b). This characteristic behavior is qualitatively consistent with those obtained by the HF approximation shown in Fig. 3. The notable difference between the ED and HF results is that spectral weights at low energies ($4.0\lesssim\varepsilon\lesssim-3.0$) in $D^{<}(\varepsilon,\tau_{\rm pr})$ decrease more largely in the ED results, which may be due to the correlation effect [51, 52].
@@
Figure 8: Momentum dependence of $A^{<}_{k}(\varepsilon,\tau_{\rm pr})$ for (a) $A_{p}=0$, (b) $A_{p}=1.2$ and $\omega_{p}=1.0$, (c) $A_{p}=2.4$ and $\omega_{p}=1.0$, and (d) $A_{p}=2.4$ and $\omega_{p}=5.0$. Other parameters are $\tau_{\rm pr}=10$, $\tau_{p}=3$, and $N=18$. The dashed line in each panel indicates $\varepsilon=0$. In (d), the arrows indicate the transferred spectral weights at $\varepsilon=1.8$ and $k=2\pi/9,7\pi/9$.
To examine the results in more detail, we show the momentum dependence of $A_{k}^{<}(\varepsilon,\tau_{\rm pr})$ for different $\omega_{p}$ and $A_{p}$ in Fig. 8. For $A_{p}=0$ (no photo-irradiation), a large peak at each $k$ forms the valence band of the CO ground state [Fig. 8(a)]. For $A_{p}=1.2$ and $\omega_{p}=1.0$, the in-gap weights appear near $k\simeq\pi/2$ ($4\pi/9$ and $5\pi/9$) and the valence band peaks are blurred due to photoexcitation [Fig. 8(b)]. For $A_{p}=2.4$, the in-gap weights further develop and they are connected to peaks forming a dispersion similar to the valence band, shifted to higher energy, as shown in Fig. 8(c). This is consistent with the HF results for $\omega_{p}\simeq\Omega_{c}$ in Figs. 2(a) and 2(b). For $A_{p}=2.4$ and $\omega_{p}=5.0$, spectral peaks in the ground state decrease and there is a spectral transfer to $\varepsilon=1.8$ at $k=2\pi/9$, $7\pi/9$ without notable dispersion around them, as shown in Fig. 8(d). This indicates the interband excitations as in the HF results shown in Fig. 2(d).
@@
Finally, we discuss the decrease in $\phi$ due to photoexcitation by varying $\omega_{p}$ and $A_{p}$ [31, 32]. In Fig. 9, we show the two-dimensional plot of $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane for $\tau_{p}=3$ and $N=22$, where $\phi_{0}$ and $\phi_{\rm av}$ are the CO order parameter in the ground state and the time-averaged $\phi$ in the time domain of $10<\tau<30$, respectively. The structure of the two-dimensional plot can be understood essentially from the behavior of $\alpha(\omega)$ shown in Fig. 5(a); for $\omega_{p}$ with large $\alpha(\omega)$, $\phi$ efficiently decreases with small $A_{p}$, whereas for $\omega_{p}$ with small $\alpha(\omega)$, a large $A_{p}$ is needed to suppress $\phi$. In Appendix B, we show the two-dimensional plot with longer pulse width of $\tau_{p}=10$ for which the correspondence with $\alpha(\omega)$ is more apparent. In the HF results [31, 32], the decrease in $\phi$ is basically small when $\omega_{p}$ is not resonant with the collective mode: we present the corresponding two-dimensional plot within the HF theory in Appendix C. However, this does not hold in the ED results, which may be due to photoinduced collective motions of CO with nonzero momenta [49] that are not captured in the HF approximation.
@@ Acknowledgements.
This work was supported by JSPS KAKENHI Grant Numbers 23K03309, 23K25826, 23K03333, and 25H00838.
Appendix A Linear optical absorption and optical conductivity
In this appendix, we compare our results of the linear optical absorption spectrum $\alpha(\omega)$ shown in Figs. 1(a) and 5(a) with optical conductivity spectra calculated within the linear response theory. Considering that the power absorption $P$ due to continuous electric field of light is estimated as $P\propto\sigma_{\rm reg}(\omega)E_{\rm cw}^{2}$ where $\sigma_{\rm reg}(\omega)$ and $E_{\rm cw}$ are the regular part of optical conductivity and the amplitude of the light electric field, respectively, $\alpha(\omega)$ is given as $\alpha(\omega)\simeq P\Delta T\propto\sigma_{\rm reg}(\omega)(\omega A_{\rm cw})^{2}\Delta T$ with $\Delta T$ being the measurement duration for the increment in total energy. In the linear response theory, $\sigma_{\rm reg}(\omega)$ is calculated as
$\sigma_{\rm reg}(\omega)=-\frac{1}{N\omega}\langle\Psi_{0}|j\frac{1}{\omega-\mathcal{H}+E_{0}+i\eta}j|\Psi_{0}\rangle,$ (15)
where $|\Psi_{0}\rangle$ is the ground-state wave function, $E_{0}$ the ground state energy, and $j$ the current operator that is defined as
$j=-it\sum_{i}(c^{\dagger}_{i}c_{i+1}-c^{\dagger}_{i+1}c_{i}).$ (16)
We first discuss the results obtained by the HF theory. In Fig. 10(a), we show $\omega^{2}\sigma_{\rm reg}(\omega)$ for $V=3$ and $\eta=0.1$ where we depict $\alpha(\omega)$ in Fig. 1(a) for comparison. The quantity $\omega^{2}\sigma_{\rm reg}(\omega)$ has a single peak at $\omega=\Delta_{\rm CO}$ corresponding to the CO gap in the ground state accompanied by a continuum-like tail above it, which is in contrast to $\alpha(\omega)$ that has a peak at $\omega=\Omega_{c}\sim\Delta_{\rm CO}/2$ due to the collective mode and the continuum above the CO gap. This discrepancy comes from the fact that in $\sigma_{\rm reg}(\omega)$ the HF order parameters are treated as static quantities, whereas $\alpha(\omega)$ calculated by the time-dependent Schrödinger equation takes account of their dynamics at the level of the random phase approximation by which the collective mode can be captured.
To see this point explicitly, we define $\alpha_{\rm RB}(\omega)$ as $\alpha(\omega)$ for a rigid band system where the HF order parameters are artificially fixed during the time evolution under the electric field in Eq. (13). We show the result of $\alpha_{\rm RB}(\omega)$ in Fig. 10(a) indicating that it has only a single peak at $\omega=\Delta_{\rm CO}$ and has no peak at $\omega=\Omega_{c}$, which is the same spectral feature as that of $\omega^{2}\sigma_{\rm reg}(\omega)$.
Next, we examine the results with the ED method. In Fig. 10(b), we show $\omega^{2}\sigma_{\rm reg}(\omega)$ for $V=3$ and $\eta=0.1$ with $N=18$ and $22$ together with the corresponding results of $\alpha(\omega)$ presented in Fig. 5(a). It is apparent that, apart from an overall scale factor, the spectral feature of $\omega^{2}\sigma_{\rm reg}(\omega)$ almost coincides with that of $\alpha(\omega)$. This demonstrates that $\omega^{2}\sigma_{\rm reg}(\omega)$ calculated exactly by using the many-body wave function essentially gives the linear absorption spectrum, which is in contrast to the case of the HF theory.
Figure 10: (a) $\omega^{2}\sigma_{\rm reg}(\omega)$ calculated within the HF theory, which is compared with $\alpha(\omega)$ shown in FIg. 1(a). We also show $\alpha_{\rm RB}(\omega)$ that indicates $\alpha(\omega)$ for the rigid band system. We note that $\omega^{2}\sigma_{\rm reg}(\omega)$ and $\alpha_{\rm RB}(\omega)$ are depicted in a linear scale. (b) $\omega^{2}\sigma_{\rm reg}(\omega)$ obtained by the ED method for $N=18$ and $22$, which are compared with $\alpha(\omega)$ shown in Fig. 5(a). We use $V=3$ and $\eta=0.1$.
Appendix B ED results for the decrease in $\phi$ on $\{\omega_{p},A_{p}\}$ plane with $\tau_{p}=10$
Figure 11: Two-dimensional plot of $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane for $\tau_{p}=10$ obtained by many-body dynamics with the ED. We use $V=3$ and $N=22$.
In this Appendix, we discuss the effects of the pulse width $\tau_{p}$ on the decrease in $\phi$ due to photoexcitation. In Fig. 11, we show the two-dimensional plot of $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane for $\tau_{p}=10$ and $N=22$ obtained by the ED method. As we have discussed in Sec. III B, the structure of Fig. 11 essentially reflects that of $\alpha(\omega)$ presented in Fig. 5(a). Compared to the case with $\tau_{p}=3$ shown in Fig. 9, each frequency range over which $\phi$ largely decreases becomes narrower. This is because the light with longer pulse width excites the system over a narrower frequency window. In particular, a large decrease in $\phi$ seen for the low-$\omega_{p}$ region ($\omega_{p}\simeq 0.5$) in Fig. 9 comes from high frequency components contained in the short pulsed light with $\tau_{p}=3$ and is not due to an adiabatic response, which is indicated by the results with $\tau_{p}=10$ (Fig. 11) where no noticeable decrease in $\phi$ appears in that region.
Appendix C HF results for the decrease in $\phi$ on $\{\omega_{p},A_{p}\}$ plane
Figure 12: Two-dimensional plot of $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane obtained by the time-dependent HF method. We use $V=3$, $\tau_{p}=3$, and $N=400$.
In Fig. 12, we show $(\phi_{0}-|\phi_{\rm av}|)/\phi_{0}$ on the $\{\omega_{p},A_{p}\}$ plane obtained by the time-dependent HF method with $N=400$ where the other parameters are the same as those of the ED results shown in Fig. 9. When $\omega_{p}$ is resonant to the collective mode ($\omega_{p}=\Omega_{c}\simeq\Delta_{\rm CO}/2$), a large decrease in $\phi$ appears, whereas a relatively weak response appears around $\omega_{p}=\Delta_{\rm CO}$ where the interband excitations suppress the CO. This result is qualitatively the same as those obtained in our previous HF studies with different parameter sets [31, 32] and can be understood from the frequency dependence of $\alpha(\omega)$ shown in Fig. 1(a).